Please solve with steps! Section 11 General Equilibrium Continued 1. Suppose there are two consumers, A and B, and two goods, X and Y. Consumer As ut
Answer #1

Answer * Given collect from the data * let us assume that the data as followed * Now we have to apome that the data of followNow X=2Y = 3 l then zu and 34 simillary B has 3 y =* Budget constrain is Px.x+Py.y = u » { xf11 (4) = u - y = 4 Now X= (-) (W